An ice skater is spinning at 2.5 revolutions per second and has a moment of inertia of 0.85 kg m2.

Estimate her rotational angular momentum, assuming for simplicity that she can be approximated as a rigid, axially-symmetric body.
Answer #1

Here,

angular speed, wf = 2.5 rev/s

wf = 2.5 * 2pi rad/s

wf = 15.71 rad/s

moment of inertia, I = 0.85 Kg.m^2

rotational angular momentum = I * wf

rotational angular momentum = 0.85 * 15.71

rotational angular momentum = 13.34 Kg.m^2/s

the rotational angular momentum is 13.34 Kg.m^2/s